OpenStudy (anonymous):

Which are identities? Thanks you guys sooo much! I think #1, #3, and #4 are identities.

OpenStudy (apoorvk):

Right-O!!

OpenStudy (campbell_st):

start with sin^2(x) + cos^2(x) = 1 Call it equation A tan(x) = sin(x)/cos(x) for Question 1 tan(2x) = 2tan(x)/(1 - tan^2x) take the reciporcal of each term 1/2tan(x) = (1-tan^2(x))/(2tan(x)) 2. Expand the perfect square (sin(x) -cos(x))(sin(x)-cos(x)) = sin^2(x) - 2sin(x)cos(x) + cos^2(x) = 1 - 2sin(x)cos(x) 2sin(x)cos(x) =sin(2x) = 1 -sin(2x) might be a typo in the question 3. \[\sin(8x) = \sin(2\times 4x) = 2\sin(4x)\cos(4x)\] hope this helps